# What is the nature of Little Red Dots and what is not, MIRI SMILES edition

### v1

@@ 2.3 Redshifts
A total of 18 galaxies out of the 31 in the whole sample (55%) are counted as having spectroscopic redshifts provided by CANDELS (Guo et al., 2013), FRESCO (Oesch et al., 2023), and JADES (Bunker et al., 2023; Eisenstein et al., 2023b), all of them flagged as highly secure.
Photometric redshifts for all galaxies were estimated with eazy (Brammer et al., 2008), using either direct flux measurements for all bands or with a modified version of the program that penalizes solutions implying fluxes above upper limits. We used v1.3 templates, which include a dusty galaxy and a high-EW emission-line galaxy spectrum, and we added a new extreme emission-line galaxy template as well as a AGNtorus model (based on the spectrum in Killi et al. 2023). Based on the comparison with spectroscopic values, we chose the maximum photometric redshift as our main solution. We estimated redshifts both using only NIRCam fluxes and also adding MIRI measurements, obtaining better results when using NIRCam only (a conclusion based on the comparison with spectroscopic values).
@@
The redshift histogram of our sample is given in Figure 1. The sample has a relatively flat redshift distribution between and , peaking at . This is probably related to the fact that strong emission lines entering the F444W filter affect the selection significantly (as also seen in extreme emission line galaxies included in Pérez-González et al. 2023a). These lines are mainly [OIII]+H and H+[NII]+[SII], which translate to and for the lines lying within the F444W passband. Summarizing, the typical redshift (median and quartiles) of LRDs selected down to F444W28 mag is .
@@ 2.4 Additional SED constraints
Apart from the NIRCam and MIRI fluxes, we used Hubble ACS fluxes measured in Hubble Legacy Field images (Illingworth et al., 2016; Whitaker et al., 2019). Except for one source, not covered by NIRCam SW channels, we did not use WFC3 data. We checked that none of the LRDs in this paper are detected by Spitzer with the MIPS instrument or by Herschel with PACS and SPIRE. Taking into consideration the catalogs used in Barro et al. (2019), we imposed in our SED analysis 5 upper limits of 30 Jy for the MIPS 24 m band, and 4, 5, 9, 11, and 11 mJy for Herschel bands at 100, 160, 250, 350, and 500 m, respectively.
@@ 3.1 Average spectral energy distributions
Figure 3 indicates that the stacked SED of the SMILES LRDs agrees well with the results from previous works based on NIRCam data: the rest-frame UV up to 0.4 m nicely matches the emission of an unobscured, or slightly reddened, QSO, which is consistent with the low-luminosity scattered light component of the accretion disk. Remarkably, the average SED presents a bump around the MgII emission feature, also supporting the interpretation that the AGN has a non-negligible, even dominating, contribution to the rest-frame UV spectral range. We note also that scatter around the median suggests that there might be sources other than an AGN contributing to the UV, e.g., emission from the host galaxy.
@@
The rest-frame optical presents a very steep slope from 0.4 to 1 m consistent with the emission of a heavily obscured supermassive black hole, here indicated with the Polletta et al. (2007) torus template. In addition, aided by the inclusion of the JEMS medium-band photometry, the stack exhibits statistically significant peaks at the wavelengths where we could expect strong H+[NII]+[SII] and [OIII]+H emission. The average LRD SED might be identified with a torus up to m, also considering that even though the torus does not have optical emission lines, the lines masquerade as a steeper (dustier) continuum. In summary, up to the reddest bands covered by NIRCam (F444W, F460M, and F480M filters, typically probing up to H), the average UV-optical SED presented in this paper roughly matches a QSO spectrum plus a torus.
@@
However, the interpretation of LRD SEDs being dominated by a pure and/or complex (with a gray attenuation law) AGN model is not supported when we consider the rest-frame near-infrared (NIR) spectral region probed by MIRI, the key addition in this paper, as first presented in Williams et al. (2023a). The MIRI data (colored circles in Figure 3) and average SED are consistently lower than the predictions of the Polletta et al. (2007) torus template. A more general evaluation can be made by applying different reddening levels to a standard unreddened AGN template. In the 0.5–3 m region where the MIRI measurements fall, there is general agreement on the shape of the intrinsic template, as determined by averaging the behavior of large samples (e.g., Elvis et al., 1994; Richards et al., 2006; Glikman et al., 2006; Lyu et al., 2017). Lyu & Rieke (2022b) show that this is an appropriate average behavior including the variations such as warm dust deficient and hot dust deficient cases (Lyu et al., 2017). It is therefore the appropriate foundation for fitting the LRD photometry, but with the large attenuations required to match the SED slope. To be specific, we adopt a SED from Glikman et al. (2006), extended to the mid-infrared (MIR) with models from Siebenmorgen et al. (2015). Figure 3 shows that there is no suitable solution; if reddening is selected to match the behavior between 0.5 and 1 m, the SED falls substantially higher than the MIRI photometry near 2 m (rest). Reducing the reddening to solve this problem yields a SED too low near 1 m. Overall, the MIRI photometry indicates a change to a significantly bluer slope at 1 m and even stabilizes at a roughly flat value at 1-2 m, where stellar emission is expected to peak. This is a fundamentally different shape from AGN SEDs. There are also hints of another steepening at m, but this behavior is traced only by a small subset of LRDs detected beyond 1.6 m rest-frame (see discussion in § 5.2), deeper MIRI data at the longest wavelengths would be necessary to probe this region.
@@
The comparison of photometry and templates in Figure 3 strongly suggests that the rest-NIR emission of the LRDs (around 1-2 m) is very different from the hot-dust (T1500 K) dominated emission of the typical QSO templates used in recent LRD papers dominated by the fit of NIRCam data alone (e.g., Barro et al. 2023, Labbé et al. 2023b, or Greene et al. 2023). Reproducing the observed MIRI colors would require a more flexible modeling of the torus emission that can vary its IR luminosity (i.e., the relative flux level with respect to other components such as stars or the accretion disk) and the wavelength at which the emission peaks. This can be accomplished using either radiative transfer models (e.g., Nenkova et al., 2008; Stalevski et al., 2012, 2016; Siebenmorgen et al., 2015), modified blackbody templates (Kim et al. 2015, Hernán-Caballero et al. 2016 or Killi et al. 2023), or empirical templates that can account for separate contributions of the torus dust emission at different temperatures (e.g., warm and hot, as in Lyu et al. 2017). However, such models require significant departures from typical AGN behavior in the red and near infrared. Alternatively, the result could indicate that the SED at 0.5–2.0 m is not AGN-dominated but rather stellar-dominated, as recently discussed in Williams et al. (2023a). We demonstrate this point by plotting in Figure 3 the torus template normalized to the average SED level of LRDs at 2 m (dashed red line). If the emission at rest-frame wavelengths longer than 2 m is dominated by a torus, then the flux at around 1 m should be dominated by a different component, such as the accretion disk or, more likely, stars, whose emission peaks exactly in that spectral zone (except for very young ages).
@@
The right panel of Figure 3 shows the average SED of LRDs that are not detected by MIRI. The UV range is bluer than in the MIRI-detected case, more in line with the slope of the average QSO spectrum. The change in slope at 0.4 m is also clear (this being one of the main selection criteria of our sample). Strong emission lines might also be present, indeed we observe a similar SED rise in the [OIII] region, and 2 NIRCam points maybe revealing a strong H emission, but the small number of sources does not allow to detect emission lines in the average SED as well as in the case for MIRI detections. For these sources, the MIRI upper limits also imply that the SED flattens at m, i.e., stellar emission could dominate this spectral range. However, the SED possibly has a steepening compatible with the presence of strong hot dust emission at wavelengths around 2 m and redward, the currently available MIRI data cannot constrain this spectral region appropriately.
@@ 3.2 MIRI colors
Figure 4 illustrates further what the MIRI observations reveal about the nature of LRDs, probing further into the redshift effects. Two of the top panels show the expected NIRCam-MIRI colors for different types of templates as a function of redshift. The nature of the LRD emission in the spectral range probed by the F444W-F560W or F444W-F770W colors cannot be disentangled, both AGN and stellar dominated emission would present very similar color differences of about 0.5 mag, and the influence of emission lines in one and/or the other filter at specific redshifts complicates the problem. Only when observing at longer wavelengths with MIRI, 10–15 m, would the data start to differentiate between the models. For our detections at these wavelengths, the MIRI data reveal flat SEDs, F770W-F1000W, and F770W-F1280W values being bluer than 0.5 mag for 60% of the sample, which is consistent with stellar dominated emission presenting a 1.6 m bump. We note however that stellar emission alone leads to flat or even negative MIRI colors, whereas the median colors of the LRDs, including the upper limits, show small, but increasingly redder colors with wavelength which suggest that there is emission from another source, such as nebular continuum or dust, but not as dominant as in the QSO templates (see also discussion in § 5.2).
@@
In summary, as found in the preceding section, this more detailed analysis shows how SEDs constructed up to the reddest NIRCam bands are compatible with a dominant AGN with a composite emission: blue in the UV, with emission lines coming from the broad-line region, and starting to rise as expected for dust torus emission up to 5 m. MIRI data at 5.6 m deviates very little from this scenario, but at 7.7 m starts to tell a different story. Even longer wavelength data clearly point to a restricted contribution of a dust torus at rest-frame wavelengths around 1–2 m, where the stellar emission is supposed to peak.
@@
Both in Figures 3 and 4, we show one of the radiative transfer models of dust-rich compact nuclear starbursts and luminous infrared galaxies (LIRGs) presented in Siebenmorgen & Krügel (2007). In particular, we plot (in blue) the template for an extreme starburst with a 350 pc star-forming region with a sub-LIRG-like luminosity ( L; but LIRG-like luminosity density), with 90% of its luminosity arising from OB stars embedded in very thick [ mag] and dense ( cm hydrogen density) birth clouds. Remarkably, this stellar-only model, mainly based on the presence of highly embedded OB stars within a more general stellar population, nicely reproduces the main characteristics of LRDs across the whole spectrum: the change in spectral slope at rest-frame m and the flattening of the SED probed by MIRI. It, however, presents a steeper slope in the FUV, but we remark the stellar emission is constant across all these models and does not take into account nebular emission. The redshift dependence of the color of this template, shown in Figure 4, is also completely consistent with the measurements.
@@ 4 SED modeling codes for LRDs
In this section, we present the detailed SED fitting methods that will be applied to the analysis of individual objects in our sample to infer their physical properties. The results will be presented in the next section, concentrating on those galaxies that have spectroscopic redshifts and are detected in several MIRI bands. First, we describe here our fitting procedures. The SEDs of all our LRDs were fitted with 4 codes, each one with different assumptions and approaches: synthesizer-AGN, prospector-SF, prospector-AGN, and prospector-AGN.
@@ 4.1 The synthesizer-AGN code
For the synthesizer code (Pérez-González et al., 2003, 2008) we assumed a composite stellar population with a young and a more evolved starburst. Both star-forming events are characterized by a Star Formation History (SFH) described by delayed exponential function with timescales between 1 Myr and 1 Gyr, and with ages from 1 Myr up to the age of the Universe at the redshift of the source. The dust effect on the stellar and nebular emission is described by the Calzetti et al. (2000) law, with A(V) values ranging from 0 to 4 mag for each population, considered to have completely independent attenuations. The stellar emission is described by the Bruzual & Charlot (2003) models, assuming a Chabrier (2003) IMF with stellar mass limits between 0.1 and 100 . Nebular emission is also considered (see Pérez-González et al. 2003 for details). Apart from the 4 parameters of each stellar population, i.e., age, timescale, attenuation and metallicity, we fit the stellar mass ratio between the youngest stars and the total mass, the burst strength. In total, this implies 9 parameters to fit. Apart from the stellar population, we include in synthesizer an AGN component described by a QSO template constructed with the average spectra of QSO presented in Vanden Berk et al. (2001), and Glikman et al. (2006), extended to the far-IR with the dust torus template in Siebenmorgen et al. (2015). This model was shown in Figure 3. We add to the 9 stellar parameters to be fitted, 2 more that describe the scaling of the AGN template with respect to the stars and its attenuation. The attenuation of the AGN model is also assumed to follow the Calzetti et al. (2000) law, and we impose an energy balance criterion so the energy absorbed by dust is reradiated in the mid- and far-IR, scaling up the dust torus emission. The dust emission is also modeled (beyond what the AGN might contribute) with the radiative transfer templates of nuclear starbursts presented in Siebenmorgen & Krügel (2007). We use the templates for 350 pc star-forming regions, with different shapes parametrized by the total IR luminosity (from sub-LIRG to Hyper-LIRG ranges), visual extinction of the starburst (from 2 to 144 mag), ratio to the total luminosity (from 40-90%) of the emission arising from OB stars embedded in dense molecular clouds of a variety of hydrogen number densities (from to cm, assuming a gas-to-dust ratio of 150).
We remark that synthesizer-AGN assumes independent extinctions for the 2 stellar populations in the host (as well as for the AGN). This is different from what is done by the next codes, where the attenuation for the stellar emission is governed by common parameters (visual attenuation and attenuation law slope), although internally young stars are treated differently to old stars.

### v2

@@ 2.3 Redshifts
A total of 18 galaxies out of the 31 in the whole sample (55%) have spectroscopic redshifts provided by CANDELS (Guo et al., 2013), FRESCO (Oesch et al., 2023) data (measured by the JADES team) and JADES NIRSpec data (Bunker et al., 2023; Eisenstein et al., 2023b), all of them flagged as highly secure.
Photometric redshifts for all galaxies were estimated with eazy (Brammer et al., 2008), using either direct flux measurements for all bands or with a modified version of the program that penalizes solutions implying fluxes above upper limits. We used v1.3 templates, which include a dusty galaxy and a high-EW emission-line galaxy spectrum, and we added a new extreme emission-line galaxy template as well as an AGNtorus model (based on the spectrum in Killi et al. 2023). Based on the comparison with spectroscopic values, we chose the maximum photometric redshift as our main solution. We estimated redshifts both using only NIRCam fluxes and also adding MIRI measurements, obtaining better results when using NIRCam only (a conclusion based on the comparison with spectroscopic values).
@@
A histogram of the redshift distribution of our sample is given in Figure 1. The distribution is relatively flat between and , peaking at . This behavior is probably related to the fact that strong emission lines entering the F444W filter affect the selection significantly (as also seen in extreme emission line galaxies included in Pérez-González et al. 2023a). These lines are mainly [OIII]+H and H+[NII]+[SII], which translate to and for the lines lying within the F444W passband. Summarizing, the typical redshift (median and quartiles) of LRDs selected down to F444W28 mag is .
@@ 2.4 Additional SED constraints
Apart from the NIRCam and MIRI fluxes, we used Hubble ACS fluxes measured in Hubble Legacy Field images (Illingworth et al., 2016; Whitaker et al., 2019). Except for one source, not covered by the NIRCam SW channels, we did not use WFC3 data. We checked that none of the LRDs in this paper are detected by Spitzer with the MIPS instrument or by Herschel with PACS and SPIRE. Taking into consideration the catalogs used in Barro et al. (2019), we imposed in our SED analysis 5 upper limits of 30 Jy for the MIPS 24 m band, and 4, 5, 9, 11, and 11 mJy for Herschel bands at 100, 160, 250, 350, and 500 m, respectively.
@@ 3.1 Average spectral energy distributions
Figure 3 indicates that the stacked SED of the SMILES LRDs agrees well with the results from previous works based on NIRCam data: the rest-frame UV up to 0.4 m nicely matches the emission of an unobscured, or slightly reddened, QSO, which is consistent with the low-luminosity scattered light component of the accretion disk. Remarkably, the average SED presents a bump around the MgII emission feature, indicating that the AGN makes a non-negligible contribution to the rest-frame UV spectral range. However, the scatter around the median suggests that there might be sources other than an AGN contributing to the UV, e.g., emission from stars in the host galaxy.
@@
The rest-frame optical presents a very steep slope from 0.4 to 1 m consistent with the emission of a heavily obscured supermassive black hole, here indicated with the Polletta et al. (2007) circumnuclear torus template. In addition, aided by the inclusion of the JEMS medium-band photometry, the stack exhibits statistically significant peaks at the wavelengths where we would expect strong H+[NII]+[SII] and [OIII]+H emission. The average LRD SED might be identified with a torus up to m; although the torus does not have optical emission lines, the lines from the central engine can masquerade as a steeper (dustier) continuum. In summary, up to the reddest bands covered by NIRCam (F444W, F460M, and F480M filters, typically probing up to H), the average UV-optical SED presented in this paper roughly matches an obscured QSO spectrum plus a torus.
@@
However, the interpretation of LRD SEDs being dominated by a pure and/or complex (with a gray attenuation law) AGN model is not supported when we consider the rest-frame near-infrared (NIR) spectral region probed by MIRI, the key addition in this paper, as first presented in Williams et al. (2023a). The MIRI data (colored circles in Figure 3) and average SED are consistently lower than the predictions of the Polletta et al. (2007) torus template. A more general evaluation can be made by applying different reddening levels to a standard unreddened AGN template. In the 0.5–3 m region where the MIRI measurements fall, there is general agreement on the shape of the intrinsic template, as determined by averaging the behavior of large samples (e.g., Elvis et al., 1994; Richards et al., 2006; Glikman et al., 2006; Lyu et al., 2017). Lyu & Rieke (2022b) show that this is an appropriate average behavior including the variations such as warm dust deficient and hot dust deficient cases (Lyu et al., 2017). It is therefore the appropriate foundation for fitting the LRD photometry, but with the large attenuations required to match the SED slope. To be specific, we adopt a SED from Glikman et al. (2006), extended to the mid-infrared (MIR) with models from Siebenmorgen et al. (2015). Figure 3 shows that there is no suitable solution; if reddening is selected to match the behavior between 0.5 and 1 m, the SED falls substantially higher than the MIRI photometry near 2 m (rest). Reducing the reddening to solve this problem yields a SED too low near 1 m. Overall, the MIRI photometry indicates a change to a significantly bluer slope at 1 m and even stabilizes at a roughly flat value at 1-2 m, where stellar emission is expected to peak. This is a fundamentally different shape from AGN SEDs. There are also hints of another steepening at m, but this behavior is traced only by a small subset of LRDs detected beyond 1.6 m rest-frame (see discussion in § 5.2); deeper MIRI data at the longest wavelengths would be necessary to probe this region.
@@
The comparison of photometry and templates in Figure 3 strongly suggests that the rest-NIR emission of the LRDs (around 1-2 m) is very different from the hot-dust (T1500 K) dominated emission of the typical QSO templates used in recent LRD papers dominated by the fit of NIRCam data alone (e.g., Barro et al. 2023, Labbé et al. 2023b, or Greene et al. 2023). Reproducing the observed MIRI colors would require a more flexible modeling of the torus emission that can vary its IR luminosity (i.e., the relative flux level with respect to other components such as stars or the accretion disk) and the wavelength at which the emission peaks. This can be accomplished using either radiative transfer models (e.g., Nenkova et al., 2008; Stalevski et al., 2012, 2016; Siebenmorgen et al., 2015), modified blackbody templates (Kim et al. 2015, Hernán-Caballero et al. 2016 or Killi et al. 2023), or empirical templates that can account for separate contributions of the torus dust emission at different temperatures (e.g., warm and hot, as in Lyu et al. 2017). However, such models require significant departures from typical AGN behavior in the red and near infrared. Alternatively, the result could indicate that the SED at 0.5–2.0 m is not AGN-dominated but rather stellar-dominated, as recently discussed in Williams et al. (2023a). We demonstrate this point by plotting in Figure 3 the torus template normalized to the average SED level of LRDs at 2 m (dashed red line). If the emission at rest-frame wavelengths longer than 2 m is dominated by a torus, then the flux at around 1 m must be dominated by a different component, such as the accretion disk or, more likely, stars, whose emission peaks exactly in that spectral zone (except for very young ages).
@@
The right panel of Figure 3 shows the average SED of LRDs that are not detected by MIRI. The UV range is bluer than in the MIRI-detected case, more in line with the slope of the average QSO spectrum. The change in slope at 0.4 m is also clear (this being one of the main selection criteria of our sample). Strong emission lines might also be present; indeed we observe a similar SED rise in the [OIII] region, and two NIRCam points may be revealing strong H emission. However, the small number of sources does not support detecting emission lines in the average SED as well as in the case for MIRI detections. For these sources, the MIRI upper limits also imply that the SED flattens at m, i.e., stellar emission could dominate this spectral range. However, the SED possibly has a steepening compatible with the presence of strong hot dust emission at wavelengths around 2 m and redward; the currently available MIRI data cannot constrain this spectral region appropriately.
@@ 3.2 MIRI colors
Figure 4 illustrates further what the MIRI observations reveal about the nature of LRDs, probing further into the redshift effects. Two of the top panels show the expected NIRCam-MIRI colors for different types of templates as a function of redshift. The nature of the LRD emission in the spectral range probed by the F444W-F560W or F444W-F770W colors cannot be disentangled: both AGN and stellar dominated emission would present very similar color differences of about 0.5 mag, and the influence of emission lines in one and/or the other filter at specific redshifts complicates the problem. Only when observing at longer wavelengths with MIRI, 10–15 m, would the data start to differentiate between the models. For our detections at these wavelengths, the MIRI data reveal flat SEDs, F770W-F1000W, and F770W-F1280W values, bluer than 0.5 mag for 60% of the sample and consistent with stellar dominated emission presenting a 1.6 m bump. We note however that stellar emission alone leads to flat or even negative MIRI colors, whereas the median colors of the LRDs, including the upper limits, show small, but increasingly redder colors with wavelength which suggest that there is emission from another source, such as nebular continuum or dust, but not as dominant as in the QSO templates (see also discussion in § 5.2).
@@
In summary, as found in the preceding section, this more detailed analysis shows how SEDs constructed up to the reddest NIRCam bands are compatible with a dominant AGN with composite emission: blue in the UV, with emission lines coming from the broad-line region, and starting to rise as expected for dust torus emission up to 5 m. MIRI data at 5.6 m deviates very little from this scenario, but at 7.7 m starts to tell a different story. Even longer wavelength data clearly point to a restricted contribution of a dust torus at rest-frame wavelengths around 1–2 m, where the stellar emission is expected to peak.
@@
Both in Figures 3 and 4, we show one of the radiative transfer models of dust-rich compact nuclear starbursts and luminous infrared galaxies (LIRGs) presented in Siebenmorgen & Krügel (2007). In particular, we plot (in blue) the template for an extreme starburst with a 350 pc star-forming region with a sub-LIRG-like luminosity ( L⊙; but LIRG-like luminosity density), with 90% of its luminosity arising from OB stars embedded in very thick [ mag] and dense ( cm-3 hydrogen density) birth clouds. Remarkably, this stellar-only model, mainly based on the presence of highly embedded OB stars within a more general stellar population, nicely reproduces the main characteristics of LRDs across the whole spectrum: the change in spectral slope at rest-frame m and the flattening of the SED probed by MIRI. It, however, presents a steeper slope in the FUV, but we remark the stellar emission is constant across all these models and does not take into account nebular emission. The redshift dependence of the color of this template, shown in Figure 4, is also completely consistent with the measurements.
@@ 4 SED modeling codes for LRDs
In this section, we present the detailed SED fitting methods that will be applied to the analysis of individual objects in our sample to infer their physical properties. The results will be presented in the next section, concentrating on those galaxies that have spectroscopic redshifts and are detected in several MIRI bands. The SEDs of all our LRDs were fitted with 4 codes, each one with different assumptions and approaches: synthesizer-AGN, prospector-SF, prospector-AGN, and prospector-AGN.
@@ 4.1 The synthesizer-AGN code
For the synthesizer-AGN code (Pérez-González et al., 2003, 2008) we assumed a composite stellar population with both a young and a more evolved starburst. The stellar emission is described by the Bruzual & Charlot (2003) models, assuming a Chabrier (2003) IMF with stellar mass limits between 0.1 and 100 . Both star-forming events are characterized by a Star Formation History (SFH) described by delayed exponential function with timescales between 1 Myr and 1 Gyr, and with ages from 1 Myr up to the age of the Universe at the redshift of the source. Nebular emission is also considered (see Pérez-González et al. 2003 for details). The attenuations of the stellar and nebular emission are described by the Calzetti et al. (2000) law, with A(V) values ranging from 0 to 4 mag for each population, considered to have completely independent attenuations. The dust emission is also modeled (beyond what the AGN might contribute) with the radiative transfer templates of nuclear starbursts presented in Siebenmorgen & Krügel (2007). We use the templates for 350 pc diameter star-forming regions, with different SED shapes parametrized by the total IR luminosity (from sub-LIRG to Hyper-LIRG ranges), visual extinction of the starburst (from 2 to 144 mag), ratio to the total luminosity (from 40-90%) of the emission arising from OB stars embedded in dense molecular clouds of a variety of hydrogen number densities (from to cm3, assuming a gas-to-dust ratio of 150).
Apart from the stellar population, we include an AGN component described by a QSO template constructed with the average spectrum of QSOs presented in Vanden Berk et al. (2001) and Glikman et al. (2006), extended to the far-IR with the dust torus template in Siebenmorgen et al. (2015). This model was shown in Figure 3. The attenuation of the AGN model is also assumed to follow the Calzetti et al. (2000) law, and we impose an energy balance criterion so the energy absorbed by dust is reradiated in the mid- and far-IR, scaling up the dust torus emission.
There are 9 free parameters to describe the stellar population. Apart from the 4 parameters for each stellar population, i.e., age, timescale, attenuation and metallicity, we fit the stellar mass ratio between the youngest stars and the total mass, the burst strength. We add to the 9 stellar parameters, 2 more that describe the scaling of the AGN template with respect to the stars and its attenuation.
We remark that synthesizer-AGN assumes independent extinctions for the 2 stellar populations in the host (as well as for the AGN). This is different from what is done by the next codes, where the attenuation for the stellar emission is governed by common parameters (visual attenuation and attenuation law slope), although internally young stars are treated differently from old stars.